OpenStudy (anonymous):

How much would $300 invested at 4% interest compounded monthly be worth after 8 years? Round your answer to the nearest cent.

Nnesha (nnesha):

\[\huge\rm A= P ( 1 + \frac{ r }{ t })^{n \times t}\] formula where P= principal amount (initial ) r= interest rate (change percent to decimal ) n = compounded monthly t= time or number of years replace variables by their values and then solve :-)